# SAM-LAD: Segment Anything Model Meets Zero-Shot Logic Anomaly Detection

### v2

@@ III-C Image-level Nearest Neighbor Search
III-D Object Feature Map Generator
III-D1 FeatUp Operation
For a given query image $I_{q}$, we now possess $k$ pairs of feature maps $(f_{q},f_{r}^{i})_{i\in[1,k]}$. These deep feature maps already capture the semantics of the images. However, these feature maps lack spatial resolution, making them unsuitable for directly performing subsequent dense anomaly detection and segmentation tasks, as the model would aggressively pool information over large areas. Therefore, inspired by FeatUp[18], we have employed its pre-trained feed-forward JBU upsampler[18] to restore the spatial information lost in the existing deep feature maps while still retaining the original semantics. Considering the balance of computational resources, inference speed, and ultimate detection accuracy, we opt to upsample the original feature maps by 8$\times$ (further analysis in Section IV-F). Subsequently, we upsample each pair of feature maps to obtain the upsampled feature maps, which size is $(C\times 8H^{{}^{\prime}}\times 8W^{{}^{\prime}})$:
@@ III-D1 FeatUp Operation
III-D2 Object mask
We use SAM to generate object masks for the input image pairs $(I_{q},I_{r}^{i})_{i\in[1,k]}$. However, in actual applications, due to SAM’s superior segmentation performance, it often segments both overly detailed objects and those not intended for detection. For example, in Fig. 3(a), in the breakfast box category, a target object might be a mix of banana chips and almonds, but SAM segments individual banana chips and almonds(Fig. 3(b)), which is not what we intended. Thus, after obtaining the segmentation masks from SAM, it is necessary to further filter them to acquire the desired key object masks.
@@ III-D2 Object mask
Specifically, upon processing the image data through the SAM, we can obtain the area of each segmented object. By setting thresholds for the minimum and maximum area, we can effectively filter the segmented objects, thus isolating the key objects of interest. For each category of the scene, SAM filters are set to determine the minimum and maximum area thresholds, resulting in object masks. The key object of interest as illustrated in Fig. 4.
@@
After segmentation by SAM and filtration, we obtain the $k$ pairs of segmented object masks for the image pairs $(I_{q},I_{r}^{i})_{i\in[1,k]}$. For the segmented object mask from anomaly-free reference image $I_{r}$, the number of object classes within its object mask is invariably constant. For instance the category of the breakfast box, an anomaly-free image always contains two tangerines, one peach, a portion of cereals, and a mix of banana chips and almonds, in addition to the main breakfast box itself, totaling six objects. Another category of screw bag contains exactly two washers, two nuts, one longscrew, two screw heads, and one short screw, totaling eight objects. Hence, for the segmentation of object masks from the anomaly-free reference image $I_{r}$, we partition it into $N$ individual object masks. (For detailed $N$ values for each category, refer to the red numerical annotations of object masks in Fig. 4). For the segmented object masks from $I_{q}$, due to the presence of missing or additional objects, the number of object classes M in the segmented object masks varies. We partition it into M object masks. Ultimately, for the image pairs $(I_{q},I_{r}^{i})_{i\in[1,k]}$, we obtain $k$ pairs of M and N object masks, respectively.
III-D3 Object Feature Map
@@
III-E Object Matching
III-E1 Dynamic Channel Graph Attention
To facilitate the matching of objects between the query and reference images, we consider each object to be a keypoint and extract their respective feature map into feature descriptor vector. However, during the process of feature compression, it is inevitable that positional information will be lost. Additionally, the recalibration of channel weights, by enhancing or suppressing semantic information, renders the extraction of global information challenging. Therefore, the establishment of a Dynamic Channel Graph Attention (DCGA) module is proposed to enhance the responsiveness between objects and channels, which can explicitly capture the spatial dependencies of various objects to augment global representation. Specifically, the feature maps of all objects can be conceptualized as a graph structure. Within this concept, each individual object feature can be regarded as a vertex in the graph, and the relationships between these objects are seen as the edges. Moreover, within the feature map of each object, each individual channel can also be viewed as a vertex, with the interactions between these channels representing edges. A schematic of the DCGA mechanism is illustrated in Fig. 5.
@@ III-E1 Dynamic Channel Graph Attention
For a query image’s objects feature map $(M,C,8H^{{}^{\prime}},8W^{{}^{\prime}})$, the dimension of each object feature map is first squeezed to $C\times 1\times 1$ by a global max pooling (GMP) operation. Then, the object features in DCGA is a tensor of shape $M\times C$. Subsequently, a graph structure is employed to generate the weights for each vector.
This graph structure consists of two parts. The first part targets each object’s feature vector $(1,C)$. Specifically, two independent $C\times C$ matrices, that is, $A_{0}$ and $A_{1}$, constitute the adjacency matrices, representing the dependency relationships among channel vertices. $A_{0}$ is generated by a 1-D convolution layer and a Softmax function. It is a self-attention-based diagonal matrix designed to suppress irrelevant features. $A_{1}$ is a predefined identity matrix, representing only the vertex itself, and requires normalization. Thus, the adjacency matrix can be represented as $A=A_{0}\times A_{1}$.
@@
$S_{ij}=\frac{S_{ij}^{\prime}}{\sum_{k=1}^{M}S_{ik}},S_{ij}^{\prime}=\hat{v}_{i}\cdot\hat{v}_{j}=\sum_{d=1}^{C}\hat{v}_{id}\cdot\hat{v}_{jd}.$ (5)
Note that, when generating the adjacency matrix, the weights updating is based solely on the similarity between one object vector and others, self-similarity is set to zero, that is, the diagonal of $A_{2}$ is zero.
@@
III-E2 Object Matching Module
@@
III-F Anomaly Measurement Module
@@
To enhance the robustness of the detection results produced by the proposed SAM-LAD, we do not simply use the object masks corresponding to the $d_{i}^{Unmatched},i\in[1,Unmatched]$ as the final score map. Instead, we build an Anomaly Measurement Module (AMM) to detect anomalies for each object in the query image individually, resulting in an overall anomaly score map.
@@
$\mathcal{M}_{Non-matching}=\sum_{i=1}^{Unmatched}\mathcal{M}_{i}.$ (12)
@@
Figure 6: Architecture of the Anomaly Measurement Module (AMM).
IV Experimental verification
IV-A Datasets
TABLE I: Comprehensive Comparison Results of the Proposed SAM-LAD and Existing Methods. Training indicates whether a method requires retraining with the dataset of a new scene when detecting anomalies in that scene. ()($\uparrow$)( ↑ ) indicates that higher values represent better performance and ()($\downarrow$)( ↓ ) indicates that lower values represent better performance. Bold font indicates the best results, while underlined represents the second-best results.
@@ IV-A Datasets
SAM-LAD(ours) $\times$ Transformer $\checkmark$ $\checkmark$ 34.6 1.9 90.7 83.2 98.0 97.7
MVTec LOCO AD: The MVTec LOCO AD dataset [11] was recently released by MVTec Software GmbH, which is specifically developed for logical anomalies and comprises five object categories, each containing both structural and logical anomalies in the test set. It has a total of 2,076 anomaly-free samples and 1,568 samples for testing. Each of the 1,568 test images is either anomaly-free or contains at least one structural or logical anomaly. Specifically, logical anomalies primarily include missing elements, extra items, disordered sequences, and inappropriate combinations, among others. Pixel-level annotations are provided as the ground truth for testing.
IV-B Evaluation metrics
We use the Area Under the Receiver Operating Characteristic Curve (AUROC) score as a threshold-free metric to evaluate image-level anomaly detection. For anomaly localization in both MVTec AD and MVTec LOCO AD datasets, AUROC is also suitable for assessing structural anomalies. However, logical anomalies of the MVTec LOCO AD, e.g. a missing object, are challenging to annotate and segment on a per-pixel basis. To evaluate anomaly localization performance, we use the saturated Per-Region Overlap (sPRO) metric[11] with the per-pixel false-positive rate of 5%, which is a generalized version of the PRO metric [1]. This metric reaches saturation once the overlap with the ground truth achieves a predefined saturation threshold. All thresholds are also provided by MVTec LOCO AD dataset.
TABLE II: Quantitative detection and localization results of the SAM-LAD framework on the MVTec LOCO AD dataset. Results for each category are given as logical anomalies/structural anomalies or the average of both. Overall averages are given as logical anomalies/structural anomalies and the average of both. The results of the comparison methods are from [11], [13], and [40].
@@ IV-B Evaluation metrics
IV-C Implementation Details
@@
IV-D Comparison With the State-of-the-Art Models
TABLE III: The AUROC Results of Various Methods in MVTec AD at the Image/Pixel-level
@@ IV-D Comparison With the State-of-the-Art Models
IV-D1 Comprehensive Comparison
In the initial phase of our study, a comprehensive evaluation was conducted to compare the proposed SAM-LAD with existing methods, including AE[36], f-AnoGan[37], SPADE[38], Padim[39], Patchcore[10], THFR[14], DSKD[15], and GLCF[13]. The results are presented in Table I. Notably, all existing methods require additional training for different scenarios, resulting in limited generalization. In contrast, our method achieves zero-shot capabilities, allowing for plug-and-play functionality in any scene while ensuring optimal logical detection performance. Analyzing the network structures employed in these methods, it was observed that most existing approaches are built upon convolutional neural networks (CNNs), with only a few utilizing vision transformers, which are still in the early stages of exploration. A significant limitation of CNN-based models lies in their inability to capture global semantics effectively, which is crucial for logical anomaly detection.
@@ IV-D1 Comprehensive Comparison
In terms of inference efficiency and computational requirements, our SAM-LAD employs a transformer structure and upsampled feature maps, which affects the inference time compared to other schemes. However, the FLOPs of SAM-LAD is still within acceptable limits. When considering benchmark performance, the SAM-LAD demonstrates advanced capabilities in Mvtec AD for detecting structural anomalies and Mvtec LOCO AD for identifying logical anomalies, owing to its proficiency in robust segment and object matching.
IV-D2 MVTec LOCO AD
The comparative results are presented in Table II. Compared to the best baseline method and the SoTA method, our framework improves +25.6 and +0.7 on image-level AUROC, +45.0 and +5.4 on pixel-level sPRO, respectively.
TABLE IV: Quantitative AUROC Comparison Results on the DigitAnatomy Dataset
@@ IV-D2 MVTec LOCO AD
Fig. 10 depicts the efficacy of each method (EfficientAD-M, GCAD, SPADE[38], DRAEM[3], GLCF) on the MVTec LOCO AD dataset concerning both structural and logical anomalies. The results indicate that our proposed strategy showcases strong capabilities in identifying both structural and logical anomalies. It is noteworthy to mention that, in comparison to the current best SoTA method, the pixel-level detection performance for logical anomalies has witnessed a significant increment of 9.1 through our framework. This highlights our method’s superior proficiency in discerning logical anomalies, which aligns perfectly with our foundational intent. It enables an in-depth comprehension of the logical interrelations among each object within the entire scene and proficiently identifies the corresponding logical discrepancies.
@@
IV-D3 MVTec AD
Beyond affirming the performance of our SAM-LAD in detecting logical anomalies, we further appraise its capability to identify structural anomalies on the MVTec AD dataset, which exclusively encompasses structural anomalies. To be specific, given that each subset within the MVTec AD dataset features a scene with a single object against a regular background, we facilitate the segmentation into this single object by employing the SAM filter for subsequent matching and computation. Owing to the one-to-one fixed matching paradigm, the ultimate detection results are directly yielded by the AMM, thereby serving as an indirect assessment of the efficacy of our proposed AMM.
We compare SAM-LAD with several SoTA methods, which are AE-SSIM[36], PMB-AE[9], MKD[44], RD4AD[46], RIAD[45], DRAEM[3], Padim[39], Patchcore[10], C-FLOW AD[47], and GLCF[13]. Table III shows the quantitative comparison results. The proposed SAM-LAD achieves remarkable image-level anomaly detection results and pixel-level anomaly localization results, obtaining an AUROC of 98.0/97.7 across 15 categories. Remarkably, SAM-LAD exhibits significantly better performance compared to baseline methods. Moreover, when compared with the SoTA methods such as Patchcore and C-FLOW, our framework achieves comparable detection and localization accuracy. A selection of qualitative results on the MVTec AD dataset is depicted in Fig. 9.
In conclusion, the proposed SAM-LAD showcases exceptional performance in the context of industrial anomaly detection. Fig. 11 offers a comprehensive comparison between SAM-LAD and existing methods regarding their image-level and pixel-level detection capabilities on the MVTec AD and MVTec LOCO AD datasets. The results unequivocally indicate that our framework attains the highest precision in anomaly detection and localization across these benchmarks, thereby evidencing the efficacy and versatility of our SAM-LAD in addressing an array of anomalies, including localized structural anomalies and complex logical anomalies.
@@ IV-D3 MVTec AD
IV-D4 DigitAnatomy
To further validate the efficacy of the proposed SAM-ALD in detecting logical anomalies, we conduct a comparative experiment employing the DigitAnatomy dataset. A gamut of comparative assessments was performed, utilizing an array of methods such as AE[36], GANomaly[48], f-AnoGAN[37], and SQUID[35], coupled with the integrative Fastflow[49] and Patchcore[10], in conjunction with Draem[3], MKD[44], RD4AD[46], and GLCF[13].
The comparative results are delineated in Table IV and the qualitative results are shown in Fig. 12. The results show that our framework markedly transcends existing methods. In particular most efficacious GLCF method, our novel SAM-LAD method realizes a substantial enhancement, accruing a gain of +13.7 in AUROC. Albeit the SQUID method was devised alongside the DigitAnatomy dataset, it garners merely an unassuming AUROC of 55.7. Conversely, our SAM-LAD framework attains a formidable AUROC of 92.3, thereby accentuating its preeminent efficacy.

### v4

@@ 3.3 Image-level Nearest Neighbor Search
3.4 Object Feature Map Generator
3.4.1 FeatUp Operation
For a given query image $I_{q}$, we now possess $k$ pairs of feature maps $(f_{q},f_{r}^{i})_{i\in[1,k]}$. These deep feature maps already capture the semantics of the images. However, these feature maps lack spatial resolution, making them unsuitable for directly performing subsequent dense anomaly detection and segmentation tasks, as the model would aggressively pool information over large areas. Therefore, inspired by FeatUp[20], we have employed its pre-trained feed-forward JBU upsampler[20] to restore the spatial information lost in the existing deep feature maps while still retaining the original semantics. Considering the balance of computational resources, inference speed, and ultimate detection accuracy, we opt to upsample the original feature maps by 8$\times$ (further analysis in Section 4.6). Subsequently, we upsample each pair of feature maps to obtain the upsampled feature maps, which size is $(C\times 8H^{{}^{\prime}}\times 8W^{{}^{\prime}})$:
@@ 3.4.1 FeatUp Operation
3.4.2 Object mask
We use SAM to generate object masks for the input image pairs $(I_{q},I_{r}^{i})_{i\in[1,k]}$. However, in actual applications, due to SAM’s superior segmentation performance, it segments all potential objects in an image that are overly sensitive objects or overly generalized. For example, in Fig.3(a), in the breakfast box category, a target object might be a mix of banana chips and almonds, but SAM segments individual banana chips and almonds(Fig.3(b)), which is not what we intended. Thus, after obtaining the segmentation masks from SAM, it is necessary to filter them further to acquire the desired key object masks.
@@ 3.4.2 Object mask
Specifically, upon processing the image data through the SAM, we can obtain the area of each segmented object. By setting thresholds for the minimum and maximum area, we ignore objects that are not intended for detection and select the key objects of interest. For each scene category, SAM filters are set to determine the minimum and maximum area thresholds, resulting in object masks. The key object of interest is illustrated in Fig 4.
@@
After segmentation by SAM and filtration, we obtain the $k$ pairs of segmented object masks for the image pairs $(I_{q},I_{r}^{i})_{i\in[1,k]}$. For the segmented object mask from anomaly-free reference image $I_{r}$, the number of object classes within its object mask is invariably constant. For instance, the category of the breakfast box, an anomaly-free image always contains two tangerines, one peach, a portion of cereals, and a mix of banana chips and almonds, in addition to the main breakfast box itself, totaling six objects. Another category of screw bag contains exactly two washers, two nuts, one long screw, two screw heads, and one short screw, totaling eight objects. Hence, for the segmentation of object masks from the anomaly-free reference image $I_{r}$, we partition it into $N$ individual object masks. (For detailed $N$ values for each category, refer to the red numerical annotations of object masks in Fig.4). For the segmented object masks from $I_{q}$, due to the presence of missing or additional objects, the number of object classes M in the segmented object masks varies. We partition it into M object masks. Ultimately, for the image pairs $(I_{q},I_{r}^{i})_{i\in[1,k]}$, we obtain $k$ pairs of M and N object masks, respectively.
3.4.3 Object Feature Map
@@
3.5 Object Matching
3.5.1 Dynamic Channel Graph Attention
To facilitate the matching of objects between the query and reference images, we consider each object to be a keypoint and extract their respective feature map into a feature descriptor vector. However, during the process of feature compression, it is inevitable that positional information will be lost. Additionally, the recalibration of channel weights by enhancing or suppressing semantic information renders the extraction of global information challenging. Therefore, the establishment of a Dynamic Channel Graph Attention (DCGA) module is proposed to enhance the responsiveness between objects and channels, which can explicitly capture the spatial dependencies of various objects to augment global representation. Specifically, the feature maps of all objects can be conceptualized as a graph structure. Within this concept, each individual object feature can be regarded as a vertex in the graph, and the relationships between these objects are seen as the edges. Moreover, within the feature map of each object, each individual channel can also be viewed as a vertex, with the interactions between these channels representing edges. A schematic of the DCGA mechanism is illustrated in Fig.5.
@@ 3.5.1 Dynamic Channel Graph Attention
For a query image’s objects feature map $(M,C,8H^{{}^{\prime}},8W^{{}^{\prime}})$, the dimension of each object feature map is first squeezed to $C\times 1\times 1$ by a global max pooling (GMP) operation. Then, the object features $f_{\text{in }}$ in DCGA is a tensor of shape $M\times C$. Subsequently, a graph structure is employed to generate the weights for each vector.
This graph structure consists of two parts. The first part targets each object’s feature vector $(1,C)$. Specifically, two independent $C\times C$ matrices, that is, $A_{0}$ and $A_{1}$, constitute the adjacency matrices, representing the dependency relationships among channel vertices. $A_{0}$ is a predefined identity matrix, representing only the vertex itself, and requires normalization. $A_{1}$ is a self-attention-based diagonal matrix designed to suppress irrelevant features, which is defined as follows:
$A_{1}=\operatorname{softmax}\left(Wf_{\text{in }}\right),$ (4)
where $W$ denotes the pre-trained weight of the 1-D convolution[36]. Thus, the adjacency matrix can be represented as:
$A=A_{0}\times A_{1}.$ (5)
@@
$S_{ij}=\frac{S_{ij}^{\prime}}{\sum_{k=1}^{M}S_{ik}^{\prime}},S_{ij}^{\prime}=\hat{v}_{i}\cdot\hat{v}_{j}=\sum_{d=1}^{C}\hat{v}_{id}\cdot\hat{v}_{jd}.$ (7)
Note that when generating the adjacency matrix, self-similarity is set to zero, meaning the diagonal of $A_{2}$ is zero.
@@
3.5.2 Object Matching Module
@@
3.6 Anomaly Measurement Module
@@
OMM employs global object feature vectors for matching, making it a coarse-level semantic feature matching method suitable for logical anomaly detection. However, in real-world scenarios, structural anomalies and logical anomalies often coexist, with structural anomalies generally being subtle. In such cases, OMM demonstrates insensitivity to these minor defects, leading to erroneous matching of structurally anomalous objects with those in reference images. Consequently, relying solely on the unmatched object mask $d_{i}^{Unmatched},i\in[1,Unmatched]$ is inadequate for computing the final anomaly map. Instead, we build an Anomaly Measurement Module (AMM) to detect anomalies for each object in the query image individually, resulting in an overall anomaly score map.
@@
$\mathcal{M}_{Non-matching}=\sum_{i=1}^{Unmatched}\mathcal{M}_{i}.c$ (14)
@@
Table 1: Comprehensive Comparison Results of the Proposed SAM-LAD and Existing Methods. Training indicates whether a method requires retraining with the dataset of a new scene when detecting anomalies in that scene. ()($\uparrow$)( ↑ ) indicates that higher values represent better performance and ()($\downarrow$)( ↓ ) indicates that lower values represent better performance. Bold font indicates the best results, while underlined represents the second-best results.
@@
SAM-LAD(ours) $\times$ Transformer $\checkmark$ $\checkmark$ 54.7 8.9 90.7 83.2 98.4 98.5
4 Experimental results
Table 2: Quantitative detection and localization results of the SAM-LAD framework on the MVTec LOCO AD dataset. Results for each category are given as logical anomalies/structural anomalies or the average of both. Overall averages are given as logical anomalies/structural anomalies and the average of both. The results of the comparison methods are from [12], [15], and [43].
@@ 4 Experimental results
4.1 Datesets
MVTec LOCO AD: The MVTec LOCO AD dataset [12] was recently released by MVTec Software GmbH, which is developed explicitly for logical anomalies and comprises five object categories, each containing both structural and logical anomalies in the test set. It has a total of 2,076 anomaly-free samples and 1,568 samples for testing. Each of the 1,568 test images is either anomaly-free or contains at least one structural or logical anomaly. Specifically, In the test images containing logical anomalies, the number and types of objects are variable. In the normal setting, all objects should adhere to specific logical constraints. Logical anomalies deviate from these constraints, manifesting as missing, extra, wrong location, or inappropriate object combinations. Pixel-level annotations are provided as the ground truth for testing.
4.2 Evaluation metrics
We use the Area Under the Receiver Operating Characteristic Curve (AUROC) score as a threshold-free metric to evaluate image-level anomaly detection. For anomaly localization in both MVTec AD and MVTec LOCO AD datasets, AUROC is also suitable for assessing structural anomalies. However, logical anomalies of the MVTec LOCO AD, e.g., a missing object, are challenging to annotate and segment on a per-pixel basis. To evaluate anomaly localization performance, we use the saturated Per-Region Overlap (sPRO) metric[12] with the per-pixel false-positive rate of 5%, which is a generalized version of the PRO metric [1]. This metric reaches saturation once it overlaps with the ground truth and achieves a predefined saturation threshold. All thresholds are also provided by the MVTec LOCO AD dataset.
4.3 Implementation Details
@@
4.4 Comparison With the State-of-the-Art Models
Table 3: The AUROC Results of Various Methods in MVTec AD at the Image/Pixel-level
@@ 4.4 Comparison With the State-of-the-Art Models
4.4.1 Comprehensive Comparison
In the initial phase of our study, a comprehensive evaluation was conducted to compare the proposed SAM-LAD with existing methods, including AE[38], f-AnoGan[39], SPADE[40], Padim[41], Patchcore[11], THFR[16], DSKD[17], and GLCF[15]. The results are presented in Table 1. Notably, all existing methods require additional training for different scenarios, resulting in limited generalization. In contrast, our method achieves zero-shot capabilities, allowing plug-and-play functionality in any scene while ensuring optimal logical detection performance. Analyzing the network structures employed in these methods, it was observed that most existing approaches are built upon convolutional neural networks (CNNs), with only a few utilizing vision transformers, which are still in the early stages of exploration. A significant limitation of CNN-based models lies in their inability to capture global semantics effectively, which is crucial for logical anomaly detection.
In terms of inference efficiency and computational requirements, our SAM-LAD employs a transformer structure and upsampled feature maps, which affects the inference time compared to other schemes. However, the FLOPs of SAM-LAD are still within acceptable limits. When considering benchmark performance, the SAM-LAD demonstrates advanced capabilities in Mvtec AD for detecting structural anomalies and Mvtec LOCO AD for identifying logical anomalies, owing to its proficiency in robust segment and object matching.
Table 4: Quantitative AUROC Comparison Results on the DigitAnatomy Dataset
@@ 4.4.1 Comprehensive Comparison
4.4.2 MVTec LOCO AD
The comparative results are presented in Table 2. Compared to the best baseline method and the SoTA method, our framework improves by +25.6 and +0.7 on image-level AUROC, +45.0 and +5.4 on pixel-level sPRO, respectively.
Fig.10 depicts the efficacy of each method (EfficientAD-M, GCAD, SPADE[40], DRAEM[3], GLCF) on the MVTec LOCO AD dataset concerning both structural and logical anomalies. The results indicate that our proposed strategy showcases strong capabilities in identifying both structural and logical anomalies. It is noteworthy that, compared to the best SoTA method, the pixel-level detection performance for logical anomalies has witnessed a significant increment of 9.1 through our framework. This highlights our method’s superior proficiency in discerning logical anomalies, which aligns perfectly with our foundational intent. It enables an in-depth comprehension of the logical interrelations among each object within the entire scene and proficiently identifies the corresponding logical discrepancies.
4.4.3 MVTec AD
Beyond affirming the performance of our SAM-LAD in detecting logical anomalies, we further appraise its capability to identify structural anomalies on the MVTec AD dataset, which exclusively encompasses structural anomalies. To be specific, given that each subset within the MVTec AD dataset features a scene with a single object against a regular background, we facilitate the segmentation into this single object by employing the SAM filter for subsequent matching and computation. Owing to the one-to-one fixed matching paradigm, the ultimate detection results are directly yielded by the AMM, thereby serving as an indirect assessment of the efficacy of our proposed AMM. We compare SAM-LAD with several SoTA methods, which are AE-SSIM[38], PMB-AE[10], MKD[47], RD4AD[49], RIAD[48], DRAEM[3], Padim[41], Patchcore[11], C-FLOW AD[50], and GLCF[15]. Table 3 shows the quantitative comparison results. The proposed SAM-LAD achieves remarkable image-level anomaly detection results and pixel-level anomaly localization results, obtaining an AUROC of 98.4/98.5 across 15 categories. Remarkably, SAM-LAD exhibits significantly better performance compared to baseline methods. Moreover, compared with the SoTA methods such as Patchcore and C-FLOW, our framework achieves comparable detection and localization accuracy. A selection of qualitative results on the MVTec AD dataset is depicted in Fig.9.
In conclusion, the proposed SAM-LAD showcases exceptional performance in the context of industrial anomaly detection. Fig.11 comprehensively compares SAM-LAD and existing methods regarding their image-level and pixel-level detection capabilities on the MVTec AD and MVTec LOCO AD datasets. The results unequivocally indicate that our framework attains the highest precision in anomaly detection and localization across these benchmarks, thereby evidencing the efficacy and versatility of our SAM-LAD in addressing an array of anomalies, including localized structural anomalies and complex logical anomalies.
4.4.4 DigitAnatomy
To further validate the efficacy of the proposed SAM-ALD in detecting logical anomalies, we conduct a comparative experiment employing the DigitAnatomy dataset. Since the DigitAnatomy dataset contains only logical anomalies, we can utilize a lightweight version of SAM-LAD for its detection. Specifically, the lightweight version calculates the anomaly map in AMM exclusively for non-matched objects. Thus, the Eq.15 is modified as:
$\mathcal{M}_{final}=\mathcal{M}_{Non-matching}=\sum_{i=1}^{Unmatched}\mathcal{M}_{i}.$ (16)
A gamut of comparative assessments was performed, utilizing an array of methods such as AE[38], GANomaly[51], f-AnoGAN[39], and SQUID[42], coupled with the integrative Fastflow[52] and Patchcore[11], in conjunction with Draem[3], MKD[47], RD4AD[49], and GLCF[15]. The comparative results are delineated in Table 4 and the qualitative results are shown in Fig.12. The results show that our framework markedly transcends existing methods. In particular, the most efficacious GLCF method, our novel SAM-LAD method, realizes a substantial enhancement, accruing a gain of +13.7 in AUROC. Albeit the SQUID method was devised alongside the DigitAnatomy dataset, it garners merely an unassuming AUROC of 55.7. Conversely, our SAM-LAD framework attains a formidable AUROC of 92.3, thereby accentuating its preeminent efficacy.
c